OpenStudy (anonymous):

HW Help Find the midpoints of the following segments using MIDPOINT FORMULA (5,4) , (5,-3)

OpenStudy (jdoe0001):

same as before \(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 5}}\quad ,&{\color{blue}{ 4}})\quad &({\color{red}{ 5}}\quad ,&{\color{blue}{ -3}}) \end{array}\qquad \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2} \right)\)

OpenStudy (anonymous):

okay i tried my best to come up with 5 , 0.5 but with the 0.5 what would that be in a fractional format? @jdoe0001

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

well 10/2 = 5 or 5/1 1/2 = 0.5 or well 1/2 :)

OpenStudy (anonymous):

or would my answer be ( 5, 1/2) im not sure @jdoe0001

OpenStudy (jdoe0001):

either one is correct just a matter of which one you're expected to give one is decimal, the other rational, both correct though